OpenStudy (anonymous):

a chili recipe calls for six pounds of ground beef for 25 serving how many pounds are needed for 30 servings 30 servings

OpenStudy (fibonaccichick666):

so if I set up the proportion could you solve it?

OpenStudy (anonymous):

Alright, so listen up. There are 6 pounds of ground beef 25 servings right? In order to find how much pounds of beef for 30 servings, you need to find how many pounds of beef is in each serving. In order to get that you need to take your 6 pounds of beef and divide it by the amount of servings which is 25 like so... 6/25 = 0.24 Every serving has about 0.24 pounds of beef. So now you take that one serving of beef (0.24) and multiply that with how many servings you want to have, in this case, it is 30 servings. 0.24 * 30 = 7.2 pounds of ground beef Decimals are always nasty numbers, so think of it as a percentage out of a 100%. 0.2 is 20% out of a 100%, so to find out what 20% of a pound is, you need to go to ounces. There are 16 ounces in a pound. Finding a percentage would be like this: 0.2 * 16 = 3.2 ounces So now you have your pounds of ground beef needed for 30 servings! To serve your chili recipe 7 pounds and 3.2 ounces is needed for 30 servings. And I'm sure that going on to the minuscule amounts beyond ounces is ridiculous... Hope this helped!